OpenStudy (anonymous):

How many miles equal 26 kilometers? [1 kilometer = 0.62 mile] 2.38 miles 8.06 miles 16.12 miles 26.62 miles

OpenStudy (anonymous):

@LynFran @GeniousCreation @hamza123

OpenStudy (lynfran):

whats 26*0.62=?

OpenStudy (anonymous):

16.12

OpenStudy (anonymous):

Correct.

OpenStudy (lynfran):

1 kilometer=(1*0.62) mile 26 kilometer= (26*0.62) mile correct

OpenStudy (anonymous):

That would be your answer.

OpenStudy (anonymous):

C. 16.12

OpenStudy (anonymous):

@LynFran @GeniousCreation thx so much

OpenStudy (anonymous):

Glad to help. :)